can you solve this for me? 0.96+0.20x=0.16(x+12)
Found 2 solutions by shree840, ewatrrr:
Answer by shree840(260)

0.96+0.20x=0.16(x+12)
0.96+0.20x=0.16x+1.92
0.20x-0.16x=1.92-0.96
0.04x=0.96
x=0.96/0.04=24

Answer by ewatrrr(24785)


Hi
Solving in the usual manner regardless of the decimals
0.96+0.20x=0.16(x+12)
0.96+0.20x=0.16x+ 1.92
.04x = .96
x = 24
CHECKING our Answer***
.96 + 4.8 = 5.76 = 3.84 + 1.92